# Using the rejection sampling for finding tests

### v1

@@ 1 Introduction
Statistical hypothesis testing is a cornerstone of statistical inference. Numerous tests have been developed for various hypothesis testing problems, such as assessing statements about group means, testing for statistical independence (see, e.g., Reshef et al.,, 2018; Kuismin,, 2025), assess whether two samples came from the same underlying distribution (see, e.g., Giovanni and Franceschini,, 1987; Chen et al.,, 2025), test the normality of the data (see, e.g., Zhou and Shao,, 2014) or determining whether two categorical variables are associated, to mention just a few examples from the extensive body of statistical tests (see, e.g., Lehmann and Romano,, 2005; Kanji,, 2006).
@@
2 Materials and method
Let $X=(x_{1},\ldots,x_{p})^{\top}$ be a $p$-dimensional real random vector and let $\textbf{X}=(X_{1},\ldots,X_{n})$ denote the data matrix. Following the notation of Casella and Berger, (2024), the general framework of statistical testing can be described as follows. Let $\theta$ denote a population parameter (vector) belonging to a parameter space $\Theta$. Let $\Theta_{0}$ be some subset of $\Theta$ and $\Theta_{0}^{c}$ is the complement of $\Theta_{0}$. In this paper $H_{0}$ and $H_{A}$ are used to denote the null and the alternative hypothesis, respectively. Consider testing $H_{0}:\theta\in\Theta_{0}$ against $H_{A}:\theta\in\Theta_{0}^{c}$.
@@ 2 Materials and method
A statistical hypothesis test is defined using a test statistic $T(X_{1},\ldots,X_{n})=T(\textbf{X})$. In most cases, the test says that $H_{0}$ should be rejected if $T(\textbf{X})$ is greater than (less than) or equal a certain constant $c$. This constant $c$ is called the rejection threshold, and the rejection region can be written as: $\{\textbf{X}\mid T(\textbf{X})\geq c\}$ where $c\in\mathbb{R}$.
@@ 2.1 The accept-reject algorithm
Rejection sampling, also known as the accept-reject algorithm (hereafter AR algorithm), is one of the conceptually simplest algorithms used to generate samples from arbitrary probability distributions of any dimension (see, e.g., Robert and Casella,, 2010, pp. 51–57). Let $f$ be a function presenting the functional form of the density of interest. For simplicity, $f$ is referred as the probability density function (pdf) of the distribution of interest; the algorithm is identical in case of probability mass function (pmf). $f$ is sometimes called as the target distribution. Use $X\sim f$ to indicate that the variable $X$ follows a specific but unknown multivariate probability distribution with density $f$. Let $g$ denote a known univariate or multivariate probability distribution with known pdf ($g$ is sometimes called as the proposal distribution). Let $D>0$ be a constant such that $f(X)\leq Dg(X)$ The AR algorithm can be outlined into the following four steps:
@@
The empirical efficiency of the AR algorithm can be evaluated using the so-called probability of acceptance, which is defined as follows: Let $\text{I}(x)$ denote the indicator function of the event $x>U$, where $U\sim\text{Unif}(0,1)$. It is assumed that $N\geq 1$ rounds of rejection sampling are performed. The probability of acceptance is given as the ratio of accepted samples to the total number of simulation rounds $N$:
@@
To elaborate how the AR algorithm can be used to find tests, consider a a goodness-of-fit problem: $H_{0}:f(X_{i})=f_{0}(X_{i})$ for all $X_{i}$, $i=1,\ldots,n$ where $f$ is a population pdf and $f_{0}$ is a hypothesized pdf, e.g., multivariate normal density. Compare this to (1) and draw a parallel between $f_{0}$ and the target distribution of the AR algorithm. Under the null hypothesis, the density $g$ from which the data are drawn should agree with the density $f_{0}$. Instead of generating pseudo-random numbers from $g$, the observed samples $X_{1},\ldots,X_{n}$ are used as input to the algorithm. The likelihood ratios $f_{0}(X_{i})/g(X_{i})$, $i=1,\ldots,n$, can then be used to assess the degree of agreement between the hypothesized pdf $f_{0}$ and the data-dependent density $g$. In particular, the resulting probability of acceptance (1) should equal one if the observations are consistent with the null hypothesis ($f_{0}$). Intuitively, this corresponds the case that the target distribution $f_{0}$ and proposal distribution $g$ are equal; setting $D=1$ represents an uncommon case when using the AR algorithm but a logical choice when investigating if the data provides sufficient evidence against the specified theoretical distribution $f_{0}$.
However, $\rho$ depends on an external random variable $U$ meaning that the value of $\rho$ changes although the likelihood ratios $f(X_{i})/g(X_{i})$, $i=1,\ldots,n$ do not change. Therefore, instead of using (1) directly, the expected value of $\rho$ is used as the test statistic. The AR test statistic is defined as follows: Let $\text{I}(x)$ be the indicator function of the event $x>U$ where $U\sim Unif(0,1)$. Let $E(Y)$ be the expected value of a random variable $Y$. Let $T(\mathbf{X})=n^{-1}\sum_{i=1}^{n}\text{I}[f(X_{i})/g(X_{i})]$. The AR test statistic is the expected value of $T(\mathbf{X})$,
$\rho(\textbf{X})=E\left[T(\textbf{X})\right].$ (2)
The AR test is any test that has the rejection region $\{\textbf{X}\mid\rho(\textbf{X})\leq c\}$, where $0\leq c\leq 1$. From this point onward, notation is slightly abused: $T(\mathbf{X})$ is used to denote averages of indicator variables, each depending on differently defined likelihood ratios.
Here, two tests derived from the AR framework are considered. The first is a goodness-of-fit test, which can be defined as follows:
@@
Let $f_{0}$ be the claimed population pdf $f$ under $H_{0}:f=f_{0}$. Let $T(\mathbf{X})=n^{-1}\sum_{i=1}^{n}\text{I}[f_{0}(X_{i})/\widehat{f}(X_{i})]$ be the mean of indicator variables depending on the ratios between the theoretical density $f_{0}$ and empirical density $\widehat{f}$, evaluated at the sample values. Use the AR test statistic (2) with rejection region $\{\textbf{X}\mid\rho(\textbf{X})\leq c\}$, where $0\leq c\leq 1$ for testing $H_{0}:f=f_{0}$ against $H_{A}:f\neq f_{0}$.
@@
The AR test is not limited to these tests; with small modifications it can be used in many testing problems. For example, in Kuismin, (2025) the author used a similar test to examine if two univariate random variables are statistically independent. The definition of the AR test is flexible: in principle, it can be defined with respect to different assumptions about the proposal distribution. For instance, while testing claims about the population covariance matrix (see, e.g., Bodnar et al.,, 2025), one could use the multivariate normal distribution, inverse Wishart distribution, the Lewandowski-Kurowicka-Joe distribution (Lewandowski et al.,, 2009), or even estimate the population distribution from data for the proposal and target distribution.
@@ 2.2 The distribution of the test statistic
Computing the distribution of $T(\textbf{X})$ by simulation is straightforward: By running the AR algorithm multiple times, the resulting values of $T(\textbf{X})$ can be used to approximate its distribution with respect to the data X. This allows evaluating the evidence provided by the AR test without resampling the data. The resulting interval estimates are credible intervals with a clear probabilistic interpretation. Moreover, it can be shown that the statistic $nT(\textbf{X})$ follows a Poisson binomial distribution when the goodness-of-fit test, as defined in Definition 2, is applied.
Let $b_{i}=\text{I}[f_{0}(X_{i})/\widehat{f}(X_{i})]$, $i=1,\ldots,n$ as defined in Definition 2. From the properties of the Bernoulli distribution and the definition of the Poisson binomial distribution, it follows that
@@
where $\text{PoiBin}(\rho_{1},\ldots,\rho_{n})$ denotes a Poisson binomial distribution with parameters $\rho_{i}=P(b_{i}>U)=\text{min}\{1,b_{i}\}$, $i=1,\ldots,n$. When $\rho_{1}=\cdots=\rho_{n}=\rho^{\prime}$, the Poisson binomial distribution reduces to the binomial distribution $\text{Bin}(n,\rho^{\prime})$.
@@
The null distribution of test statistic $\rho(\textbf{X})$ must be specified in order to determine the rejection threshold $c$ and the $p$-value of the AR test. A natural choice is $c=1$, but this makes the test too liberal since $H_{0}$ would be rejected for any $\rho(\textbf{X})<1$. In addition, the null distribution is not well-defined in this case because the variance of the test statistic is zero. Thus, setting $c=1$ is not appropriate for all applications.
From the central limit theorem, it follows that the null distribution of $\rho(\textbf{X})$ is approximately normal. However, this normal approximation typically provides a crude estimate of the null distribution, since $\rho(\textbf{X})$ is close to one under $H_{0}$. For this reason, the null distribution is estimated using a simulation approach. In particular, a Monte Carlo procedure is used to determine the rejection threshold $c$ and, consequently, the $p$-value of the AR test (see, e.g., Howes,, 2023, for Monte Carlo significance testing). The following Monte Carlo significance testing procedure is used in this paper:
@@
    Repeat steps 1.- 2. $M$ times to get $\rho(\tilde{\textbf{X}})_{1},\ldots,\rho(\tilde{\textbf{X}})_{M}$ of the AR test statistic.
@@
    Use $\rho(\tilde{\textbf{X}})_{1},\ldots,\rho(\tilde{\textbf{X}})_{M}$ to estimate the null distribution of $\rho(\textbf{X})$ and calculate the Monte Carlo $p$-value, $p_{ar}=\sum_{m=1}^{M}[I(\rho(\tilde{\textbf{X}})_{m}\leq\rho(\textbf{X}))+1]/(M+1)$.
@@
In the following sections, the empirical statistical power of three distinct AR tests is assessed through a comparative Monte Carlo simulation study, where the rejection threshold $c$ is determined by Monte Carlo simulations.
@@ 2.3 Comparing equality of group means
In this first example, the difference between group means is investigated, which represents one of the fundamental problems in statistical inference dating back to the pioneering work of Student, (1908). Let $\mu=(\mu_{1},\ldots,\mu_{p})^{\top}$ denote a population mean vector of length $p\geq 2$. Consider testing:
@@
Here the AR test defined in Definition 3 is applied for testing the hypothesis described above. In this example, it is demonstrated how incorporating certain distributional assumptions can improve the computational efficiency of the AR test. In particular, it is illustrated how the AR test can be constructed using a sufficient statistic. Here, the test is constructed based on the sample mean vector, which is a sufficient statistic for $\boldsymbol{\mu}$. The AR test for the sample mean vector is defined as follows: Let $\bar{X}=[\sum_{i=1}^{n}x_{j}^{i}/n]_{p\times 1}$ be the sample mean vector of length $p$. Denote $\mu_{0}=(\mu_{1}^{0},\ldots,\mu_{p}^{0})^{\top}$ be a parameter vector of length $p$ under $H_{0}$. Examine the value of the following statistic
@@
$\rho(\textbf{X})=E\left[T(\textbf{t})>U\right]=P\left[T(\textbf{t})>U\right]=\min\{T(\textbf{t}),1\}.$ (5)
@@
The empirical power of the AR test is compared with the paired $t$-test, two-sample $t$-test, and the Likelihood Ratio (LR) test when $p=2$. According to the Neyman-Pearson Lemma, the LR test is uniformly the most powerful (UMP) test when considering a simple hypothesis $H_{0}:\mu_{d}=0$ against $H_{A}:\mu_{d}=\eta$ where $\mu_{d}=\mu_{1}-\mu_{2}$ and $\eta$ is the true difference between group means. The LR test is based on the assumption that all parameter values are known and does not involve parameter estimation. It is included here as a baseline test to illustrate the power that the optimal test could have when $p=2$ (see, e.g. Casella and Berger,, 2024, pp. 387–394). The sample size $n$ is set as 52 which is the minimal number of observations needed for the two sided paired $t$-test to have power at least $0.8$ when assuming that the group variances are equal and the effect size (Cohen’s D) is small (Cohen,, 1988).
For this simulation example, Let $X=(x_{1},x_{2})^{\top}\sim N(\mu,\Sigma)$ where $\mu=(\mu_{1},\mu_{2})^{\top}$ and $\Sigma=[\sigma_{ij}]$, $i,j\in\{1,2\}$ is a symmetric and positive definite matrix. In particular, consider $H_{0}:\mu_{1}=\mu_{2}$ against $H_{A}:\mu_{1}\neq\mu_{2}$. Here it is assumed that group variances are equal, $\sigma_{ii}=1$ and $\sigma_{ij}\in[-0.99,0.99]$. The power for each test is estimated when the difference between population mean parameters change. In particular, let $\eta=\mu_{1}-\mu_{2}$ when $\eta\in\{0.4,0.6,0.8\}$ at different values of the population correlation $\sigma_{ij}$. The empirical power of different tests is the ratio of significant tests of 10000 Monte Carlo samples at $\alpha=0.05$. Figure 2 illustrates the empirical power with respect to the population correlation.

### v2

@@ 1 Introduction
Statistical hypothesis testing is a cornerstone of statistical inference. Numerous tests have been developed for various hypothesis testing problems, such as assessing statements about group means, testing for statistical independence (see, e.g., Reshef et al., 2018; Kuismin, 2025), assess whether two samples came from the same underlying distribution (see, e.g., Giovanni and Franceschini, 1987; Chen et al., 2025), test the normality of the data (see, e.g., Zhou and Shao, 2014) or determining whether two categorical variables are associated, to mention just a few examples from the extensive body of statistical tests (see, e.g., Lehmann and Romano, 2005; Kanji, 2006).
@@
2 Proposed framework
Following Casella and Berger (2024), let $\theta$ denote a population quantity of interest (which may be a parameter vector or a more general functional) taking values in a space $\Theta$. Let $\Theta_{0}$ be some subset of $\Theta$, and let $\Theta_{0}^{c}$ denote its complement. Throughout this paper, $H_{0}$ and $H_{A}$ are used to denote the null and alternative hypotheses, respectively. In the parametric setting, the general format of null and alternative hypothesis is $H_{0}:\ \theta\in\Theta_{0}\quad\text{against}\quad H_{A}:\ \theta\in\Theta_{0}^{c}$. For example, one might be interested in testing $H_{0}:\ \theta=0\quad\text{against}\quad H_{A}:\ \theta\neq 0$ where $\theta$ is a population average parameter.
@@ 2 Proposed framework
Let $X=(x_{1},\ldots,x_{p})^{\top}$ be a $p$-dimensional real-valued random vector, and let $\mathbf{X}=(X_{1},\ldots,X_{n})$ denote the corresponding data matrix consisting of $n$ independent observations. A statistical hypothesis test is defined using a test statistic $T(X_{1},\ldots,X_{n})=T(\textbf{X})$. In most cases, the test says that $H_{0}$ should be rejected if $T(\textbf{X})$ is greater than (less than) or equal a certain constant $c$. This constant $c$ is called the rejection threshold, and the rejection region can be written as: $\{\textbf{X}\mid T(\textbf{X})\geq c\}$ where $c\in\mathbb{R}$.
@@ 2.1 The accept-reject algorithm
Rejection sampling, also known as the accept-reject algorithm (hereafter AR algorithm), is one of the conceptually simplest algorithms used to generate samples from arbitrary probability distributions of any dimension (see, e.g., Robert and Casella, 2010, pp. 51–57). Let $f$ be a function presenting the functional form of the density of interest. For simplicity, $f$ is referred as the probability density function (pdf) of the distribution of interest; the algorithm is identical in case of probability mass function (pmf). $f$ is sometimes called as the target distribution. Use $X\sim f$ to indicate that the variable $X$ follows a specific but unknown multivariate probability distribution with density $f$. Let $g$ denote a known univariate or multivariate probability distribution with known pdf ($g$ is sometimes called as the proposal distribution). Let $D>0$ be a constant such that $f(X)\leq Dg(X)$ The AR algorithm can be outlined into the following four steps:
@@
The empirical efficiency of the AR algorithm can be evaluated using the so-called probability of acceptance, which is defined as follows: Let $\text{I}(x)$ denote the indicator function/random variable of the event $x>U$, where $U\sim\text{Unif}(0,1)$. It is assumed that $N\geq 1$ rounds of rejection sampling are performed. The probability of acceptance is given as the ratio of accepted samples to the total number of simulation rounds $N$, and it can be estimated using the following estimator:
@@
2.2 Deriving a test using rejection sampling
To elaborate how the AR algorithm can be used to find tests, consider a practical goodness-of-fit problem for multivariate distributions: $H_{0}:f(X_{i})=f_{0}(X_{i})$ for all $X_{i}$, $i=1,\ldots,n$ where $f$ is a population pdf and $f_{0}$ is a hypothesized pdf, e.g., multivariate normal density. Compare this to (1) and draw a parallel between $f_{0}$ and the target distribution of the AR algorithm. Under the null hypothesis, the density $g$ from which the data are drawn should agree with the density $f_{0}$. Instead of generating pseudo-random numbers from $g$, the observed samples $X_{1},\ldots,X_{n}$ are used as input to the algorithm. The likelihood ratios $f_{0}(X_{i})/g(X_{i})$, $i=1,\ldots,n$, can then be used to assess the degree of agreement between the hypothesized pdf $f_{0}$ and density $g$. Let $g$ be a density estimate and assume, that $H_{0}$ is true. In this special case, the resulting probability of acceptance (1) should equal one if the observations are consistent with the null hypothesis ($f_{0}$). Intuitively, this corresponds the case that the target distribution $f_{0}$ and proposal distribution $g$ are equal; setting $D=1$ represents an uncommon case when using the AR algorithm but a logical choice when investigating if the data provides sufficient evidence against the specified theoretical distribution $f_{0}$.
However, $\rho$ depends on an external random variable $U$ meaning that the value of $\rho$ changes although the likelihood ratios $f(X_{i})/g(X_{i})$, $i=1,\ldots,n$ do not change. Therefore, instead of using (1) directly, the expected value of $\rho$ is used as the test statistic. The AR test statistic is defined as follows: Let $\text{I}(x)$ be the indicator function of the event $x>U$ where $U\sim Unif(0,1)$. Let $E(Y)$ be the expected value of a random variable $Y$. Let $T(\mathbf{X})=n^{-1}\sum_{i=1}^{n}\text{I}[f(X_{i})/g(X_{i})]$. The AR test statistic is the expected value of $T(\mathbf{X})$ with respect to $U$,
$\rho(\textbf{X})=E_{U}\left[T(\textbf{X})\right].$ (2)
The AR test is any test that has the rejection region $\{\textbf{X}\mid\rho(\textbf{X})\leq c\}$, where $0\leq c\leq 1$. From this point onward, notation is slightly abused: $T(\mathbf{X})$ is used to denote averages of indicator variables, each depending on differently defined likelihood ratios. In the next, two tests derived from the AR framework are considered.
2.2.1 AR goodness-of-fit-test
The first test is a goodness-of-fit test, which was already introduced in the previous section. More formally, it can be defined as follows:
@@ 2.2.1 AR goodness-of-fit-test
Let $f_{0}$ be the claimed population pdf $f$ under $H_{0}:f=f_{0}$. Let $T(\mathbf{X})=n^{-1}\sum_{i=1}^{n}\text{I}[f_{0}(X_{i})/\widehat{f}(X_{i})]$ be the mean of indicator variables depending on the ratios between the theoretical density $f_{0}$ and density estimate $\widehat{f}$, evaluated at the sample values. Use the AR test statistic (2) with rejection region $\{\textbf{X}\mid\rho(\textbf{X})\leq c\}$, where $0\leq c\leq 1$ for testing $H_{0}:f=f_{0}$ against $H_{A}:f\neq f_{0}$.
2.2.2 AR test for a parameter vector
@@
The AR test is not limited to tests (2) and (3); with small modifications it can be used in many testing problems. For example, in Kuismin (2025) the author used a similar test to examine if two univariate random variables are statistically independent. The definition of the AR test is flexible: in principle, it can be defined with respect to different assumptions about the proposal distribution. For instance, while testing claims about the population covariance matrix (see, e.g., Bodnar et al., 2025), one could use the multivariate normal distribution, inverse Wishart distribution, the Lewandowski-Kurowicka-Joe distribution (Lewandowski et al., 2009), or even estimate the population distribution from data for the proposal and target distribution.
@@ 2.3 The distribution of the test statistic
Estimating the distribution of $T(\textbf{X})$ by simulation is straightforward: By running the AR algorithm multiple times, the resulting values of $T(\textbf{X})$ can be used to approximate its distribution with respect to the data X. This allows evaluating the evidence provided by the AR test without resampling the data. The resulting interval estimates are credible intervals with a clear probabilistic interpretation. Moreover, it can be shown that the statistic $nT(\textbf{X})$ follows a Poisson binomial distribution. For simplicity, consider the case when the goodness-of-fit test, as defined in Definition 2, is applied.
Let $b_{i}=\text{I}[f_{0}(X_{i})/\widehat{f}(X_{i})]$, $i=1,\ldots,n$. From the properties of the Bernoulli distribution and the definition of the Poisson binomial distribution, it follows that
@@
where $\text{PoiBin}(\rho_{1},\ldots,\rho_{n})$ denotes a Poisson binomial distribution with parameters $\rho_{i}=P(b_{i}>U)=\text{min}\{f_{0}(X_{i})/\widehat{f}(X_{i}),1\}$, $i=1,\ldots,n$. When $\rho_{1}=\cdots=\rho_{n}=\rho^{\prime}$, the Poisson binomial distribution reduces to the binomial distribution $\text{Bin}(n,\rho^{\prime})$.
@@
The null distribution of test statistic $\rho(\textbf{X})$ must be specified in order to determine the rejection threshold $c$ and the $p$-value of the AR test. A natural choice is $c=1$, but this makes the test too liberal since $H_{0}$ would be rejected for any $\rho(\textbf{X})<1$. In addition, the null distribution is not well-defined in this case because the variance of the test statistic is zero. Thus, setting $c=1$ is not appropriate for any application.
From the central limit theorem, it follows that the null distribution of $\rho(\textbf{X})$ is approximately normal. However, this normal approximation typically provides a crude estimate of the null distribution, since $\rho(\textbf{X})$ is close to one under $H_{0}$. For this reason, the null distribution is estimated using a simulation approach. In particular, a Monte Carlo procedure is used to determine the rejection threshold $c$ and, consequently, the $p$-value of the AR test (see, e.g., Howes, 2023, for Monte Carlo significance testing). The following Monte Carlo significance testing procedure is used in this paper (where a small value of $\rho(\textbf{X})$ heuristically indicates evidence against $H_{0}$):
@@
    Repeat steps 1 to 2 $M$ times to get $\rho(\tilde{\textbf{X}})_{1},\ldots,\rho(\tilde{\textbf{X}})_{M}$ of the AR test statistic.
@@
    Use $\rho(\tilde{\textbf{X}})_{1},\ldots,\rho(\tilde{\textbf{X}})_{M}$ to estimate the null distribution of $\rho(\textbf{X})$ and calculate the Monte Carlo $p$-value, $p_{ar}=\sum_{m=1}^{M}[\text{I}(\rho(\tilde{\textbf{X}})_{m}\leq\rho(\textbf{X}))+1]/(M+1)$.
@@
In the following sections, the empirical statistical power of three distinct AR tests is assessed through a comparative Monte Carlo simulation study, where the rejection threshold $c$ is determined by Monte Carlo simulations. It will be demonstrated numerically that, when $H_{0}$ is true, the probability that the AR tests incorrectly reject $H_{0}$ does not exceed the nominal significance level. Moreover, when a real effect is present, the power of the AR tests approaches one as the sample size increases. Each new AR test is also compared with state-of-the-art statistical tests.
@@ 2.4 Comparing equality of group means
This first example illustrates the parametric use of the AR test framework. In particular, the difference between group means is investigated, which represents one of the fundamental problems in statistical inference dating back to the pioneering work of Student (1908). Let $\mu=(\mu_{1},\ldots,\mu_{p})^{\top}$ denote a population mean vector of length $p\geq 2$. Consider testing:
@@
Here the AR test defined in Definition 3 is applied for testing the hypothesis described above. In this example, it is demonstrated how incorporating certain distributional assumptions can improve the computational efficiency of the AR test. It is illustrated how the AR test can be constructed using a sufficient statistic. Here, the test is constructed based on the sample mean vector, which is a sufficient statistic for $\boldsymbol{\mu}$. The AR test for the sample mean vector is defined as follows: Let $\bar{X}=[\sum_{i=1}^{n}x_{j}^{i}/n]_{p\times 1}$ be the sample mean vector of length $p$. Denote $\mu_{0}=(\mu_{1}^{0},\ldots,\mu_{p}^{0})^{\top}$ be a parameter vector of length $p$ under $H_{0}$. Examine the value of the following statistic
@@
$\rho(\textbf{X})=E_{U}\left[T(\textbf{t})>U\right]=P\left[T(\textbf{t})>U\right]=\min\{T(\textbf{t}),1\}.$ (5)
@@
The empirical power of the AR test is compared with the paired $t$-test, two-sample $t$-test, and the Likelihood Ratio (LR) test when $p=2$. According to the Neyman-Pearson Lemma, the LR test is uniformly the most powerful (UMP) test when considering a simple hypothesis $H_{0}:\mu_{d}=0$ against $H_{A}:\mu_{d}=\eta$ where $\mu_{d}=\mu_{1}-\mu_{2}$ and $\eta$ is the true difference between group means. The LR test is based on the assumption that all parameter values are known and does not involve parameter estimation. It is included here as a baseline test to illustrate the power that the optimal test could have when $p=2$ (see, e.g. Casella and Berger, 2024, pp. 387–394). The sample size $n$ is set as 52 which is the minimal number of observations needed for the two sided paired $t$-test to have power at least $0.8$ when assuming that the group variances are equal and the effect size (Cohen’s D) is small (Cohen, 1988).
For this simulation example, Let $X=(x_{1},x_{2})^{\top}\sim N(\mu,\Sigma)$ where $\mu=(\mu_{1},\mu_{2})^{\top}$ and $\Sigma=[\sigma_{ij}]$, $i,j\in\{1,2\}$ is a symmetric and positive definite matrix. Consider the hypothesis test $H_{0}:\mu_{1}=\mu_{2}$ against $H_{A}:\mu_{1}\neq\mu_{2}$. Here it is assumed that group variances are equal, $\sigma_{ii}=1$ and $\sigma_{ij}\in[-0.99,0.99]$. The power for each test is estimated when the difference between population mean parameters change. Let $\eta=\mu_{1}-\mu_{2}$ when $\eta\in\{0.4,0.6,0.8\}$ at different values of the population correlation $\sigma_{ij}$. The empirical power of different tests is the ratio of significant tests of 10000 Monte Carlo samples at $\alpha=0.05$. Figure 2 illustrates the empirical power with respect to the population correlation.